Exercise 3: Solve the following differential equation (with initial conditions) for the three cases below..Solve the following differential equation (with initial conditions) for the three cases below (by hand!). You may use whateve
Answer #1

a) let aurillary equation corresponding to homogencouy equahan 23 +32 4 -21 20 is, 2m² +3m-2=0 si.c. f(t)=0}. 2m² + 4m-m-2=0. 200) : so. D -> G+62 - 5/3 = - - City = 8/3 - 2 I by 0 , x (t) = 1.6 et 12 - 26 e 27 ts et. As ico)-3, 2 .4-262+ = 3: → C-4c) let flt) = -4t2 let Xp= A + B t + ct² - ip = Bt at ca > X = 2C. substituting above values in giren equation, we get, - 40